OpenStudy (anonymous):

The area of a playground is 20 square yards. the length of the playground is 5 times longer than its width. Find the length and width of the playground. A) length= 1 yd, width= 20yd B) length= 2 yd, width= 10yd C) length= 10 yd, width= 2yd D) length= 20 yd, width= 1yd

OpenStudy (anonymous):

area = l*w l=5w area = \((5w)(w)\\5w^2=20\\w^2=\dfrac{20}{5}\\w^2=4\\w=\pm 2\) we discard the negative answer and we're left with w=2. plug back into the givens to find length

OpenStudy (anonymous):

So C? :)

OpenStudy (anonymous):

@♂

OpenStudy (anonymous):

looks good, looks like we didn't even have to find L because there's only one choice where w=2

OpenStudy (anonymous):

yea thank you!

OpenStudy (anonymous):

mhmmmm

OpenStudy (anonymous):

If the length is 5 times if the wide,then there is only one option in the list, we don't have to do it so complicated.